OpenStudy (anonymous):

Find the center (h, k) and radius r of the circle with the given equation 1. (x - 6)2 + (y - 5)2 = 81 2. x2 + y2 + 12x + 14y + 85 = 9 3. (x - 4)2 + y2 = 16

OpenStudy (anonymous):

(x-9)^2 + (y-7)^2 =36 find center (h,k) and find radius r of this circle

OpenStudy (anonymous):

1.C(6,5) r=9 3.C(4,0) r=4

OpenStudy (hoblos):

the general form of the equation of a circle id (x-h)^2 + ( y-k)^2 = r^2 where (h,k) is the center and r is the radius

OpenStudy (anonymous):

thanks for the explanation@hoblos...

OpenStudy (anonymous):

how do you solve it to achieve those answers

OpenStudy (anonymous):

you dont solve it compare the question with hoblos's equation and get ur answer!!!

OpenStudy (anonymous):

\[\huge (x-\underbrace{k}_{center})^2+(y-\underbrace{h}_{center})^2=\underbrace{r^2}_{radius}\]

OpenStudy (anonymous):

whoa, good one @Kreshnik

OpenStudy (anonymous):

@Sarkar thanks :)

OpenStudy (anonymous):

how do you write those @Kreshnik ???

OpenStudy (anonymous):

\huge (x- \underbrace{k}_{center} )^2+(y- \underbrace{h}_{center} )^2=\underbrace{r^2}_{radius} you just need to copy this and pasto into "Equation" button. :) and this will appear: \[\huge (x- \underbrace{k}_{center} )^2+(y- \underbrace{h}_{center} )^2=\underbrace{r^2}_{radius}\]

OpenStudy (anonymous):

okay thanks@kreshnik

OpenStudy (anonymous):

\[\huge (x- \overbrace{k}^{center} )^2+(y- \underbrace{h}_{center} )^2=\underbrace{r^2}_{radius}\] lololo.... hahaha...

OpenStudy (anonymous):

@Sarkar You're welcome :)